OpenStudy (anonymous):

What is the remainder when x²+5 is divided by x+1 ?

OpenStudy (anonymous):

let do the division and find out

OpenStudy (cwrw238):

use remainder theorem - find (f-1)

OpenStudy (anonymous):

Use remainder theorem

jimthompson5910 (jim_thompson5910):

According to the remainder theorem, if P(x) is divided by x-k and you get a remainder of r, then P(k) = r So because x^2 + 5 = (-1)^2 + 5 = 1 + 5 = 6, this means that P(-1) = 6 and that the remainder is 6

OpenStudy (anonymous):

since, x+1 is a factor, by remainder theorem the remainder is f(-1)=6